What is the rate of return on a $1,000 face-value, 4% coupon bond that was purchased for $950 and sold one year later for $920?

Group of answer choices

19%

6%

4%

2%
Answer #1

4% coupon value would be 40

thus purchasing price =950

Benefit or amount received from bond =920+40=960

Thus rate of return=10/950*100=1.05

thus choose rate of return=2%